Subject: Beamtools Update Friday Jan 31, 2003
From: James Conway <jwc@snf.stanford.edu>
Date: Fri, 31 Jan 2003 17:30:16 -0800

 
Greetings:

Previous report to this list did not get through the mail servers.
(CWIII's daily report..)
 

SEM Practicum training suspended this week and hopefully will resume FEB. 6, 2003

New SOP procedures for the SEM's will be coming out shortly.  Please stand by...

SEMHITACHI S-800 IS UP!

The Hummer sputter coater has had its target replaced with a Au-Pd target and will now render more uniform films.
This is not a deposition system for making contacts or first layers to
dep. plate up layers onto; it is certainly not for jewelry or general
purpose coating.

Unknown user(s) have been performing very long duration coating runs of
100's of microns of deposition..  The Au target had been completely
depleted and run until it perforated!  These targets last for years if
the system are used properly.

Bounty for persons whom can identify user(s) whom are abusing the
equipment or running non-SEM coating processes in the system.
 


S-4160 IS UP!

We are working on improving the digital imaging capability and usefulness of this system.
Users are reminded that this is a semi-clean contamination level system.
NO AU, non-std. metals, or GaAs in the Instrument!
This system is for advanced SEM users and you must have experience on the S-800 before you can attend SEM S-4160 Training.
Please immediately report to staff members if you experience brightness or imaging problems. Do not ever make adjustments to the FE-gun settings.

EBEAM, Hitachi H-700 EBL system: Available to all qualified Users.
Available to all users while we continue to test and process Process Control Monitors.  Placing in 'yellow light mode' over weekend to indicate your results may not be to sub 100 nm specifications.

Initial PCM pattern testing resolved lines to >250 nm on STD. PMMA. (5% 495K PMMA 350 nm thickness --lot 10-03-03)
Stitching and butting errors were observed to better than 150 nm.
Sub field issues remain, currently set at 0.5 um...
We are working to quickly address these issues.

Patterns are being run over this weekend on qualified and traceable wafers and we will hopefully run final patterns and report to you these results early Tuesday.

RAITH 150 UPDATE:  Currently testing no available to users...

This week a massive effort was made by Joe Klingfus (RAITH USA), Andreas Rampe (RAITH GmbH, Dortmund, Germany), and myself to determine root causes and develop solutions to the stitching errors we have been intermittently observing on the system.  This problem first appeared in late November and slowly has grown in magnitude and frequency. It was not initially reported by users until Fu Min Wang brought several types of small defects to my attention the first week of December.

The problem:

A very small rotation of the Write Field to write field alignment was causing the stitching errors. (~0.05 degrees in rotation)

Last week we worked on the stage and mechanical system making adjustments to sine and cosine control board of the Laser stage controls, and inspecting and cleaning the kinetic mounts to this sub-system. A number of test were run with varying results. Initial test completed last Friday rendered an excellent stitching and pattern placement result.  A longer test over last weekend totally bombed out and no single write field was observed in acceptable alignment.

This week much effort was spent taking data, measuring stitching pattern fields, and brainstorming on causalities from the observations.  We gave eliminated the ELPHY pattern generator and beam deflection system as a root cause and are no back to stage mechanical and control systems. We have observed trends that indicate that the farther away the pattern is written from the location of the Align Write field procedure the worse the problem becomes.  But it was an intermittent problems to start which made testing and evaluation very difficult.

We have developed and are testing a work around procedure through the weekend to allow students to get up and running to assist us in testing this problem that should achieve in specification results for patterns smaller than 1 X 1 mm.  We have a plan of action and are moving quickly to correct the problem; bringing in suitable replacement parts and expertise from Germany to effect a stage repair/rebuild if necessary.

RAITH Champions whom wish to help us with testing should stop buy during my office hour Monday and Tuesday -- CIS 31 8:30 - 9:30 daily. (650-725-7075)
for consideration of afternoon and evening access to the system Monday and Tuesday.  We will be working over the weekend to collect more data and test our work around procedure. Full update and stats. Tuesday AM.

Thank you for your interest in E-beam Technologies here at SNF!
 

James Conway
E-beam Technology Group
SNF
650-725-7075